# MoE-CAP: Benchmarking Cost, Accuracy and Performance of Sparse Mixture-of-Experts Systems

### v2

@@ 1 Introduction
Figure 2: Highlight of MoE system benchmarking: Memory bandwidth versus Power for AI devices. Horizontal lines show the minimum bandwidth required for MoE models with different batch size configurations (reflecting different deployment scenarios). Different boxes represent device categories suited to various applications, such as edge systems for autonomous driving and robotics, low-power computers, workstations, and data centers.
@@ 4.2 CAP Analysis Model Inputs
Here, $S$ denotes the amount of accessed memory, with $S_{\text{activated}}$ representing the memory required for activated parameters and $S_{\text{KV}}$ denoting the average memory size of the KV cache per token. The term $l_{\text{seq}}\,s_{\text{kv}}$ captures the required FLOPs to compute the attention score over the existing KV cache (i.e., $l_{\text{seq}}$ tokens), where $s_{\text{kv}}$ is the per-token KV cache size and may differ across various attention mechanisms. The constant $2$ in Equation 2 reflects the assumption that each parameter undergoes two operations—one multiplication and one addition—common in matrix multiplication. We omit element-wise operations because they can be fused into matrix operations without additional memory access, and their computational cost is orders of magnitude lower than that of matrix computations.
@@ 6 Insights and Takeaway Messages
MoE systems enable a broader range of low-power devices to perform inference. Personal machines typically operate in single-user inference environments, where experts are sparsely activated with a small batch size (usually as 1). In this case, many low-power hardware platforms can run large MoE models (with the support of offloading) or smaller versions of MoE models (such as Moonlight-16B) with reasonable latency and performance. This challenges the conventional belief that large MoE models must run on expensive data center systems, as demonstrated in Figure 2.
@@
MoE systems shall be co-optimized for applications and scenarios. We observe that the sparsity characteristics of MoE models are significantly affected by the applications (e.g., online inference, offline inference, fine-tuning and pre-training) and deployment scenarios (e.g., single-user vs. multi-user, small vs. large batch sizes). These characteristics determine system bottlenecks, which can be addressed by various hardware platforms—some offering significantly lower costs than conventional GPU-only solutions. We anticipate a surge in specialized MoE systems tailored to specific applications and scenarios.
New sparsity-driven benchmark and performance prediction methods are needed. Sparsity is not unique to MoE—it also plays a crucial role in emerging sparse KV-cache architectures (key in long-context reasoning) and multi-agent LLM scenarios. In these cases, sparsity fundamentally shapes system bottlenecks, emphasizing the need for new benchmarks and predictive models that natively integrate cost, performance, and accuracy considerations. Without such tools, sparse AI systems risk failing to deliver on their cost-saving potential and performance targets.

### v3

@@ 1 Introduction
Figure 2: Highlight of MoE system benchmarking: Memory bandwidth versus Power for AI devices. Horizontal lines indicate the minimum bandwidth required for MoE models to meet a specified decoding latency. Each model is evaluated under two scenarios: fully activated (large batch size) and minimum activated (batch size = 1). Blue dots represent peak memory bandwidth and TDP, while orange dots indicate bandwidth and TDP when offloading to DRAM is needed. Systems with dots above the horizontal line meet the latency requirement, while those below do not. Systems are clustered based on their target applications, such as edge systems (autonomous driving, robotics), low-power computers, workstations, and data centers.
@@ 4.2 CAP Analysis Model Inputs
Here, $S$ denotes the amount of accessed memory, with $S_{\text{activated}}$ representing the memory required for activated parameters and $S_{\text{KV}}$ denoting the average memory size of the KV cache per token. The term $l_{\text{seq}}\,s_{\text{kv}}$ captures the required FLOPs to compute the attention score over the existing KV cache (i.e., $l_{\text{seq}}$ tokens), where $s_{\text{kv}}$ is the per-token KV states size and may differ across various attention mechanisms. The constant $2$ in Equation 2 reflects the assumption that each parameter undergoes two operations—one multiplication and one addition—common in matrix multiplication. We omit element-wise operations because they can be fused into matrix operations without additional memory access, and their computational cost is orders of magnitude lower than that of matrix computations.
@@ 6 Insights and Takeaway Messages
MoE systems enable a broader range of devices to perform inference. Personal machines typically operate in single-user inference environments, where experts are sparsely activated with a small batch size (usually as 1). In this case, many low-power hardware platforms can run large MoE models (with the support of offloading) or smaller versions of MoE models (such as Moonlight-16B) with reasonable latency and performance. This challenges the conventional belief that large MoE models must run on expensive data center systems, as demonstrated in Figure 2.
@@
MoE systems should be co-designed to align with specific applications and deployment scenarios. We observe that the sparsity characteristics of MoE models are significantly affected by the applications (e.g., online inference, offline inference, fine-tuning and pre-training) and deployment scenarios (e.g., single-user vs. multi-user, small vs. large batch sizes). These characteristics determine system bottlenecks, which can be addressed by various hardware platforms—some offering significantly lower costs than conventional GPU-only solutions. We anticipate a surge in specialized MoE systems tailored to specific applications and scenarios.
New benchmarking and design principles are needed for emerging sparse AI systems.. Sparsity is not unique to MoE—it also plays a critical role in emerging sparse AI systems, such as those involving sparse KV-cache architectures (common in long-context reasoning) and module-based agentic LLM workflows. In these cases, sparsity directly influences system bottlenecks, highlighting the need for new system benchmark, profiling and design principles that inherently account for cost, performance, and accuracy. With these principles, a wide range of sparsity-aware system optimizations can be unlocked. Without them, sparse AI systems risk falling short of their cost-saving potential and performance goals.